OpenStudy (anonymous):

Suppose you choose a marble from a bag containing 3 red marbles, 5 white marbles, and 4 blue marbles. You return the first marble to the bag and then choose again. Find P(red and blue) Answers A 7/12 B 2/3 C 1/12 D 1

OpenStudy (anonymous):

why dont you try and type what you have tried.

OpenStudy (farmdawgnation):

JamesWolf's suggestion sounds like an excellent idea.

OpenStudy (anonymous):

P(Red U Blue) = 3/12 * 4/12 = 1/4 * 1/3 = 1/12

OpenStudy (anonymous):

or just read rushal's answer and post another question...

OpenStudy (farmdawgnation):

Rushal, regardless of whether or not you show the steps, when you interrupt someone else trying to work with the user and provide the solution, that counts as providing "just answers" - and it's not welcome here man.

OpenStudy (anonymous):

look i was just helping him around and i didnt knew that we are not allowed to give answers straight away

OpenStudy (anonymous):

i am new to this site

OpenStudy (farmdawgnation):

Well, now you do. Part of my job is to tell you when you made a mistake, you correct it, and then we both move on happier and better off for it. Now might be a good time for you to read over our Code of Conduct to help prevent situations like this in the future: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

atleast now your less new :)

OpenStudy (anonymous):

:)